Talcon Corporation has a 11% unlevered cost of equity. The company forecasts the free cash flows shown below. The cash flows are expected to grow at a constant rate of 4% rate after Year 3.

Unlevered cost of equity 11%

Growth rate after Year 3 4% Year 1. Year 2 Year 3

Free Cash Flow 450 750 805

a. Calculate the expected free cash flow for Year 4. What might cause the free cash flow to be lower?

b. Calculate the horizon value of the unlevered operations.

c. Calculate the total value of unlevered operations at Year 0.
Answer #1

FCF for year 4 = FCF Year 3*(1+growth rate)

= 805*(1+4%)

= 837.20

Horizon value = P3 = FCF4/(k-g)

= 837.20/ (11%-4%)

= 11960

Total value of unlevered operations = FCF1/(1+k)+ FCF2/(1+k)^2 +(FCF3+P3)/(1+k)^3

= 450/1.11^1+750/1.11^2+ (805+11960)/1.11^3

= 10347.78